OpenStudy (anonymous):

Solve 2(4 – b) = 2 – b – (b – 6). a. b=0 b. b=-4 c. This is an identity and is true for all real numbers. d. there is no solution.

OpenStudy (mayankdevnani):

8-2b=2-b-b+6 8-2b=2-2b+6 8-2b=8-2b 0=0

OpenStudy (mayankdevnani):

@sbanuelos ok

OpenStudy (mayankdevnani):

b=0

OpenStudy (anonymous):

this was rong.

OpenStudy (mayankdevnani):

why??

OpenStudy (mayankdevnani):

oh yaa

OpenStudy (mayankdevnani):

8-b=2-2b+6 8-b=8-2b 8-8=-2b+b -b=0 b=0

OpenStudy (mayankdevnani):

@ajprincess am i right?

OpenStudy (mayankdevnani):

@sbanuelous

OpenStudy (mayankdevnani):

plz tell me